A few years ago, Hendersen Corp. issued preferred stock paying 8% of its par value of $46. The issue is currently selling for $36. Preferred stock flotation costs are 10% of the proceeds of the sale. What is Hendersen's cost of preferred stock? Round the answer to two decimal places.
Answer #1

Annual dividend=46*8%=$3.68

Hence cost of preferred stock=Annual dividend/Current price*(1-Flotation cost)

=3.68/(36*(1-0.1))

=3.68/32.4

=11.36%(Approx).